OpenStudy (anonymous):

can someone tell me if I have solved this correctly? Simplify the following expression... sqrt(-6) * sqrt(-42) = (-126)

OpenStudy (anonymous):

\[\sqrt{?}\sqrt{-6}*\sqrt{42}= \sqrt{-6}=\sqrt{-1}*\sqrt{6}= \sqrt{-1}* \sqrt{2}*\sqrt{3} =i *\sqrt{2} *\sqrt{3} = 3i \sqrt \sqrt{-42}=\sqrt{-1}\sqrt{42} = \sqrt{-1} *\sqrt{2} * \sqrt{21} = i *\sqrt{2} *\sqrt{21} = 21i \sqrt{-6}*\sqrt{-42} = 3i \sqrt{2} * 21i \sqrt{2} = (21*3) (\sqrt{2} \sqrt{2} )(i*i)= 63*2*-1 =126 *-1= -126\]

OpenStudy (anonymous):

0r would the answer be 1/7y^2 * 21x^3 * 1/y3 =1/7y^2 *1y^3 = 1/7y^5 *21x^ 3

OpenStudy (anonymous):

\[\sqrt{-6} \times \sqrt{-42} \implies \sqrt{(-6) \times (-42)} \implies \sqrt{6 \times 42} \implies \sqrt{ 6 \times 6 \times 7 = ??}\]

OpenStudy (anonymous):

\[\sqrt{\underline{6 \times 6} \times 7} = ??\]

OpenStudy (anonymous):

252?

OpenStudy (anonymous):

You can pull out 6 there.. In square root out of two pairs you can pull out one out of the square root brackets like: \[\sqrt{\underline{x \times x} \times y} \implies x \sqrt{y}\]

OpenStudy (anonymous):

\[6\sqrt{6*7}\] ?

OpenStudy (anonymous):

Once you pulled out 6 you will not write 6 in the square root brackets...

OpenStudy (anonymous):

\[6\sqrt{7}\] ?

OpenStudy (anonymous):

In square root brackets after pulling one from the pair we delete the pair in the square root brackets: \[\sqrt{4 \times 4 \times 5} \implies 4 \sqrt{5}\]

OpenStudy (anonymous):

Yes that is right..

OpenStudy (anonymous):

thank you:)

OpenStudy (anonymous):

Welcome dear..